given that f(7)=-10 find the value of a in the eauation f(x)=a(x-4)^2+8
Answer by ewatrrr(24785)


Hi
given that f(7)= -10 find the value of a
f(x)=a(x-4)^2+8
-10 = a(3)^2 + 8
-18/9 = a
-2 = a